OpenStudy (anonymous):

Which inequalities are true? Choose all answers that are correct. –3 < –3.5 12 ≥ –12.5 7.5 ≤ 7.5 23 > 23

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

lets work these from last Is 23 greater than 23 ?

OpenStudy (anonymous):

NO!

OpenStudy (anonymous):

Woah srry for the caps

ganeshie8 (ganeshie8):

so the last option is wrong skip that

ganeshie8 (ganeshie8):

7.5 ≤ 7.5 how about this ? Is 7.5 less than or equal to 7.5 ?

OpenStudy (anonymous):

okay the ≤ what is that?

ganeshie8 (ganeshie8):

\(\le \) is read as "less than or equal to" it means "at most"

OpenStudy (anonymous):

They are the same so skip?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Umm xD where did eriwan goooo?

ganeshie8 (ganeshie8):

\(7.5 \le 7.5\) suppose your height is \(7.5\) and your brother height is also \(7.5\) is it okay to say, you're "at most" as tall as your brother ?

OpenStudy (anonymous):

so its a choice i would say

ganeshie8 (ganeshie8):

Yes \(7.5\le 7.5\) is a true statement so it has to be chosen

OpenStudy (anonymous):

Okay my next guess is the first one is correct!

ganeshie8 (ganeshie8):

Is having $3.5 as debt any better than having $3 as debt ?

ganeshie8 (ganeshie8):

which situation do you think is better, $3.5 debt or $3 debt ?

OpenStudy (anonymous):

Well 3.5 is 3 1/2 right?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

thats better than -3

OpenStudy (anonymous):

So its i was right @ganeshie8